# The light burden of memory: constraining primordial black holes with high-energy neutrinos

### v1

@@ Abstract
PBHs are metastable objects and emit fundamental particles as they evaporate through Hawking radiation. In the standard scenario, only PBHs with masses above $10^{15}$ grams would evaporate in a timescale longer than the current Hubble time, and could thus be considered viable dark matter candidates. However, strong constraints have been placed, allowing for PBHs as the sole dark matter component only in the asteroid mass range of $10^{17}{\rm g}\lesssim M_{\rm PBH}\lesssim 10^{22}{\rm g}$ Carr et al. (2016); Green and Kavanagh (2021); Carr et al. (2021); Carr and Kuhnel (2022). PBHs with $M_{\rm PBH}\lesssim 10^{15}$ g –which would have entirely evaporated by the time the Universe reaches its current age– may still play important roles in the production of the baryon asymmetry of the Universe Fujita et al. (2014); Hamada and Iso (2017); Morrison et al. (2019); Chen et al. (2020); Perez-Gonzalez and Turner (2021); Datta et al. (2021); Hooper and Krnjaic (2021); Jyoti Das et al. (2021); De Luca et al. (2021); Bernal et al. (2022a); Calabrese et al. (2023, 2024); Schmitz and Xu (2024); Barman et al. (2024a), gravitational wave emission Papanikolaou et al. (2021); Domènech et al. (2021); Papanikolaou (2022); Ireland et al. (2023); Domènech and Tränkle (2024), or dark matter generation Bernal and Zapata (2021a, b, c); Cheek et al. (2022a, b); Samanta and Urban (2022); Bernal et al. (2021a, b, 2022b), depending on their mass.
@@
The semi-classical phenomenon of black hole evaporation relies on the assumption that the black hole remains classical throughout its lifetime Hawking (1975). However, it is increasingly clear that this model may not be self-consistent, indicating the need for new physics, particularly regarding the information loss paradox Preskill (1992). Indeed, Hawking’s result ignores the back-reaction of emission on the quantum state of the black hole. This effect becomes however crucial when the energy of the emitted quanta is comparable to the black hole’s total energy. Recent studies Dvali (2018); Dvali et al. (2020, 2024) have suggested that the back-reaction may lead to a universal “memory burden” effect. This is caused by the fact that the information stored in a system resists its decay, due to the response of the quantum modes associated to the enthropic degrees of freedom. Hence, as the black hole’s mass falls below a certain threshold, back-reaction becomes significant, slowing evaporation and potentially extending its lifetime. This implies that black holes below $10^{15}$ grams could be still evaporating by now, thus leading to interesting phenomenological implications Barman et al. (2024b); Bhaumik et al. (2024); Barman et al. (2024c); Kohri et al. (2024). Moreover, the memory burden effect also reopens the possibility for light PBHs with masses below $10^{9}$ grams to be viable dark matter candidates Dvali et al. (2020); Alexandre et al. (2024); Thoss et al. (2024); Haque et al. (2024).
@@
In this Letter, we point out that a crucial probe of the existence of memory-burdened PBHs is represented by the observation of high-energy neutrinos as the only unimpeded messenger at high energies. Previous studies have so far explored the PBH emission of low-energy neutrinos  Halzen et al. (1995); Dasgupta et al. (2020); Lunardini and Perez-Gonzalez (2020); Calabrese et al. (2022); De Romeri et al. (2021); Bernal et al. (2022c); Liu and Ng (2024) and the existence of relic neutrinos from fully evaporated light PBHs Bugaev and Konishchev (2002a, b); Wu and Xu (2024), relying on the semi-classical approximation of PBH evaporation. We here address the possibility of detecting the emission of high-energy neutrinos from a population of PBHs with masses $M_{\rm PBH}\lesssim 10^{9}~{}{\rm g}$ evaporating today thanks to the memory burden effect. We compute the expected neutrino flux at Earth, assuming that these PBHs account for a significant fraction of the dark matter energy density (see Fig. 1). Then, by comparing our predictions with the latest and projected observations from high-energy neutrino telescopes, we establish tighter constraints on the parameter space of PBHs and the associated memory burden effect (see Fig. 2).
@@
Fig. 1 shows the all-flavor-sum neutrino flux $\Phi_{\nu}$ computed for three benchmark choices of the parameters $k$ and $M_{\rm PBH}$, taking $f_{\rm PBH}=10^{-6}$. As expected, the lighter the PBH mass, the higher the energies of the neutrinos emitted via evaporation. In all the cases, we find that the neutrino flux is dominated by the galactic component, with the primary (secondary) emission manly contributing to the peak (tail). On the other hand, the extragalactic component emission is almost negligible. In the plot, we also show the current IceCube observations given by the 7.5-year High-Energy Starting Events (HESE) data sample Abbasi et al. (2021) (black data points on the left) and the 7-year Extremely High-Energy (EHE) upper bound $E^{2}\Phi_{\nu}\leq 5.9\times 10^{-9}~{}{\rm GeV/cm^{2}/s/sr}$ from $1.0\times 10^{6}$ to $4.0\times 10^{9}$ GeV at 90% C.L. Aartsen et al. (2016) (gray data point in the middle), as well as the 3-year future sensitivity of IceCube-Gen2 Aartsen et al. (2019, 2021) (thin solid black line) and GRAND200k Álvarez-Muñiz et al. (2020) (thin dashed gray line) telescopes. We note that the scenario of memory-burden PBHs could be also probed with the currently-operating KM3NeT telescope with sensitivity to neutrinos up to few tens of PeV Adrian-Martinez et al. (2016); Aiello et al. (2024).

### v2

@@ Abstract
PBHs are metastable objects and emit fundamental particles as they evaporate through Hawking radiation. In the standard scenario, only PBHs with masses above $10^{15}$ grams would evaporate in a timescale longer than the current Hubble time, and could thus be considered viable dark matter candidates. However, strong constraints have been placed, allowing for PBHs as the sole dark matter component only in the asteroid mass range of $10^{17}{\rm g}\lesssim M_{\rm PBH}\lesssim 10^{22}{\rm g}$ Carr et al. (2016); Green and Kavanagh (2021); Carr et al. (2021); Carr and Kuhnel (2022). PBHs with $M_{\rm PBH}\lesssim 10^{15}$ g –which would have entirely evaporated by the time the Universe reaches its current age– may still play important roles in the production of the baryon asymmetry of the Universe Fujita et al. (2014); Hamada and Iso (2017); Morrison et al. (2019); Chen et al. (2020); Perez-Gonzalez and Turner (2021); Datta et al. (2021); Hooper and Krnjaic (2021); Jyoti Das et al. (2021); De Luca et al. (2021); Bernal et al. (2022a); Calabrese et al. (2023, 2024); Schmitz and Xu (2024); Barman et al. (2024a); Gunn et al. (2024), gravitational wave emission Papanikolaou et al. (2021); Domènech et al. (2021); Papanikolaou (2022); Ireland et al. (2023); Domènech and Tränkle (2024), or dark matter generation Bernal and Zapata (2021a); Gondolo et al. (2020); Bernal and Zapata (2021b, c); Cheek et al. (2022a, b); Samanta and Urban (2022); Bernal et al. (2021a, b); Sandick et al. (2021); Bernal et al. (2022b); Cheek et al. (2023); Gehrman et al. (2024); Bertuzzo et al. (2024), depending on their mass.
@@
The semi-classical phenomenon of black hole evaporation relies on the assumption that the black hole remains classical throughout its lifetime Hawking (1975). However, it is increasingly clear that this model may not be self-consistent, indicating the need for new physics, particularly regarding the information loss paradox Preskill (1992). Indeed, Hawking’s result ignores the back-reaction of emission on the quantum state of the black hole. This effect becomes however crucial when the energy of the emitted quanta is comparable to the black hole’s total energy. Recent studies Dvali (2018); Dvali et al. (2020, 2024) have suggested that the back-reaction may lead to a universal “memory burden” effect. This is caused by the fact that the information stored in a system resists its decay, due to the response of the quantum modes associated to the enthropic degrees of freedom. Hence, as the black hole’s mass falls below a certain threshold, back-reaction becomes significant, slowing evaporation and potentially extending its lifetime. This implies that black holes below $10^{15}$ grams could be still evaporating by now, thus leading to interesting phenomenological implications Balaji et al. (2024); Barman et al. (2024b); Bhaumik et al. (2024); Barman et al. (2024c); Kohri et al. (2024); Jiang et al. (2024). Moreover, the memory burden effect also reopens the possibility for light PBHs with masses below $10^{9}$ grams to be viable dark matter candidates Dvali et al. (2020, 2021); Alexandre et al. (2024); Thoss et al. (2024); Haque et al. (2024).
@@
In this Letter, we point out that a crucial probe of the existence of memory-burdened PBHs is represented by the observation of high-energy neutrinos as the only unimpeded messenger at high energies. Previous studies have so far explored the steady PBH emission of low-energy neutrinos  Halzen et al. (1995); Dasgupta et al. (2020); Lunardini and Perez-Gonzalez (2020); Calabrese et al. (2022); De Romeri et al. (2021); Bernal et al. (2022c); Liu and Ng (2024), the final neutrino burst from a nearby PBH Capanema et al. (2021); Perez-Gonzalez (2023); De Romeri et al. (2024), and the existence of relic neutrinos from fully evaporated light PBHs Bugaev and Konishchev (2002a, b); Wu and Xu (2024), relying on the semi-classical approximation of PBH evaporation. We here address the possibility of detecting the emission of high-energy neutrinos from a population of PBHs with masses $M_{\rm PBH}\lesssim 10^{9}~{}{\rm g}$ evaporating today thanks to the memory burden effect. We compute the expected neutrino flux at Earth, assuming that these PBHs account for a significant fraction of the dark matter energy density (see Fig. 1). Then, by comparing our predictions with the latest and projected observations from high-energy neutrino telescopes, we establish tighter constraints on the parameter space of PBHs and the associated memory burden effect (see Fig. 2).
@@
Fig. 1 shows the all-flavor-sum neutrino flux $\Phi_{\nu}$ computed for three benchmark choices of the parameters $k$ and $M_{\rm PBH}$, taking $f_{\rm PBH}=10^{-6}$. As expected, the lighter the PBH mass, the higher the energies of the neutrinos emitted via evaporation. In all the cases, we find that the neutrino flux is dominated by the galactic component, with the primary (secondary) emission manly contributing to the peak (tail). On the other hand, the extragalactic component emission is almost negligible. Similar neutrino fluxes may also arise from the mergers of galactic memory-burdened PBHs that have resumed Hawking evaporation Zantedeschi and Visinelli (2024). In the plot, we also show the current IceCube observations given by the 7.5-year High-Energy Starting Events (HESE) data sample Abbasi et al. (2021) (black data points on the left) and the 7-year Extremely High-Energy (EHE) upper bound $E^{2}\Phi_{\nu}\leq 5.9\times 10^{-9}~{}{\rm GeV/cm^{2}/s/sr}$ from $1.0\times 10^{6}$ to $4.0\times 10^{9}$ GeV at 90% C.L. Aartsen et al. (2016) (gray data point in the middle), as well as the 3-year future sensitivity of IceCube-Gen2 Aartsen et al. (2019, 2021) (thin solid black line) and GRAND200k Álvarez-Muñiz et al. (2020) (thin dashed gray line) telescopes. We note that the scenario of memory-burden PBHs could be also probed with the currently-operating KM3NeT telescope with sensitivity to neutrinos up to few tens of PeV Adrian-Martinez et al. (2016); Aiello et al. (2024).